### v2

@@ Methods
First, we provide an overview of the experimental set-up for guiding ESM3 backbone structure generation with fold class labels from CATH [26], a hierarchical protein structure classification scheme. This enables both high-level specification of overall secondary structure compositions and more fine-grained descriptions about particular folds. Concretely, we trained a fold classifier, $p_{\phi}(\mathbf{c}\,|\,\mathbf{x})$ that predict the CATH labels, $\mathbf{c}$, from the discrete structure tokens outputted by the structure encoder of ESM3, $\mathbf{x}$, on the set of CATH-labeled protein structures from the Protein Data Bank (PDB). We combined this fold class predictor with the structure generation track of ESM3, $p_{\theta}(\mathbf{x})$, to obtain a fold class conditioned backbone structure generative model, $p_{\theta,\phi}(\mathbf{x}\,|\,\mathbf{c})$.
Fold classifier
@@
Now we describe additional details of the fold class predictor. Since ESM3 operates on discrete structure tokens, we need a predictor which predicts the fold class label of a protein from its structure token representation. To this end, we first trained a fold class predictor that takes in discrete structure tokens as inputs. The model architecture consists of an encoder-only Transformer trunk. The final per-position embedding is mean-pooled, then fed into three parallel linear layers, each projecting the embedding to a vector with the dimension equal to the number of classes for each of the three levels of CATH hierarchies: C (“Class”), A (“Architecture”) and T (“Topology/Fold”). The C level describes the overall secondary structure-content of a structure (e.g., “mainly $\alpha$”, “mainly $\beta$”, and “mixed $\alpha/\beta$), the A level roughly clusters structures by the overall arrangement of secondary structure in space (e.g.“beta barrel”, “roll”), and the T level assigns structures by how the secondary structure elements are connected and arranged (e.g.“immunoglobulin-like”, “TIM barrel”) [26]. The model was trained on the PDB dataset following the same pre-processing procedure as in Geffner et al. [18] with a total of 214,564 structures, categorized into 5 C classes, 43 A classes, and 1,336 T classes. We fed each structure through the structure encoder of ESM3 to obtain its structure token representation. The dataset is randomly divided into training, validation, and test sets at a ratio of 8:1:1, ensuring that at least one protein from each class is included in the test set whenever possible.
@@
The model was trained with a standard cross-entropy loss applied to each of the CATH level. Some proteins have multiple CATH labels due to having multiple domains and CATH being a domain-level annotation. For such proteins, we followed Geffner et al. [18] and randomly sampled one domain label as the ground truth for each training iteration, encouraging the model to predict equal probabilities for all domain labels of a protein. During evaluation, if the model predicts any of the correct labels, the prediction is considered correct. For each protein, the loss is calculated for each level of its available CATH label. The transformer trunk of the model has a hidden dimension size of 512, 16 attention heads and 5 encoder layers. In total, the model has around 19 million parameters. The model is trained using the AdamW optimizer [72, 73] with a learning rate of 0.0001, weight decay coefficient of 0.01, a dropout rate of 0.2, a batch size of 128 and a gradient accumulation step of 4. The model is trained for around 26,000 training steps on a single NVIDIA RTX 6000 GPU. We did not systematically tune the hyperparameters of the model. On its test set, the model achieves high accuracies for each level of CATH based on the metrics of Micro Accuracy and Macro Accuracy (Table A[74].
@@
The fold classifier used for guidance needs to take in partially masked inputs. We initialized the model weights of this “noisy” classifier with those of the unnoised classifier, and trained it on the same training dataset as the unnoised classifier. We trained the noisy classifier for around 28,000 additional training steps, with the same hyperparameters as the unnoised classifier.
Fold class guidance
@@
We performed fold class guidance with ESM3 across the three hierarchies of CATH, following the evaluation set-up of Geffner et al. [18]. We sampled backbone structures conditioning on different classes across three levels of the CATH hierarchies. We drew 100 backbone structure samples for each C-level class (mainly $\alpha$, mainly $\beta$, mixed $\alpha/\beta$), 30 samples for each A-level class, and 30 samples for each of the 30 randomly selected T-level classes, using either the unguided structure generation track of ESM3, $p_{\theta}(\mathbf{x})$, or ESM3 with fold class guidance, $p_{\theta,\phi}(\mathbf{x}\,|\,\mathbf{c})$. For each class, for every sample, we sampled its length uniformly among the lengths present for the proteins of that class in the AFDB, with a minimum length of 50 and maximum length of 275. For unguided generation, we used the default sampling setting of ESM3 with a sampling temperature of 0.7 and $L$ sampling steps, where $L$ is the length of the protein sequence. We also experimented with using a lower sampling temperature and found that it results in ESM3 producing predominantly mainly $\alpha$ proteins. For guided generation, we used a sampling temperature of 0.3, Euler integration with a step size $\Delta t=0.005$, sampling stochasticity $\eta=20$ and guidance strength $\gamma=100$. Additional results with different settings of sampling temperatures are in the Supplementary Information (Fig. 5).
We evaluated the generated structures on two criteria: (i) whether the fold classifier, $p_{\phi}(\mathbf{c}\,|\,\mathbf{x})$, predicts the desired class $\mathbf{c}$ as the most likely class for the generated structure $\mathbf{x}$, and (ii) whether the structures are considered designable, where designability is calculated following the procedure in Yim et al. [75]: for each generated backbone, eight sequences are sampled using ProteinMPNN [1] with a temperature of 0.1, then ESMFold [44] is used to predicted the structure of each generated sequence. The RMSD is calculated between the generated structure and the predicted structure, and a generated backbone is considered designable if the lowest RMSD—referred to as the self-consistency RMSD (scRMSD)—among the eight sampled sequences is $\leq 2$Å. For the purpose of computing aggregated metrics across multiple classes, we grouped the classes for each level by their frequency of occurrence in AFDB [76] and calculated the classification accuracy and the fraction of designable proteins within each group (Table 5). The frequency groups are defined in the same way as in Geffner et al. [18].
Detailed results
On the whole, we observed that across all levels and groups, guidance steers generation towards the desired classes (Fig. 2c). Generally, we observed that the classification accuracy is lower for the A and T level and especially so for the rarer classes. Considering that the A and T level fold class labels are more fine-grained and the constraints are harder to satisfy, and increasingly so for the rarer classes given the imbalanced nature of the fold class distribution, this result is expected. We also observed that among the samples that are classified correctly, only a moderate fraction is designable. This aligns with previous observations that ESM3 tends to produce backbones that are less designable, possibly due to the composition of its training set [18]. We also note that increasing designability does not necessarily correlate with a better capability in generating functional backbone structures, as it has been observed that many functional proteins, e.g., those in the CATH databases, are not designable by the standard definition, and backbone generation models with high designability can systematically miss parts of the structure space that are enriched for these functional proteins [77]. Since guided generation depends crucially on the capability of the underlying unguided model, we can expect structure generation to be improved as better structure generative models are developed. Examples of the generated backbone structures by guiding ESM3 with some example of CATH classes can be found in the Supplementary Information (Fig. 7). The shown examples are both classified correctly to the conditioning fold class by the fold classifier and are designable (scRMSD $\leq 2$Å).
In addition to the metrics of classification accuracy and designability, we also characterized the diversity among the generated structures. We followed Geffner et al. [18] in quantifying diversity by computing the pairwise TM-scores [78] among designable samples stratified by length. TM-scores range from zero to one, where higher scores indicate greater similarity and lower scores means higher diversity among the generated samples. Since we intended to evaluate the quality of the conditional samples, we further filtered the samples to those that are classified correctly to the conditioning fold class by the fold classifier. Since the unguided model rarely sampled structures of the desired class for the A and T level, we only evaluated the diversity of the generated samples when conditioning with classes at the C level. For each C level classes (mainly $\alpha$, mainly $\beta$, mixed $\alpha/\beta$), we sampled 10,000 structures either with unguided ESM3, or guided with ProteinGuide, and filtered the generated structures to those that are both classified to the desired class by the fold classifier and are designable. We further stratified these samples by length and computed the pairwise TM-scores among the samples for each combination of class and length. The diversity of the generated structures are comparable between the unguided model and that model with ProteinGuide (Fig. 6).
@@
Guiding base editor for improved editing activity
Overview
@@
ABE Library cloning
ABE libraries contained a diversified region of 86 amino acids (position 82-167) that was synthetically generated as an oligo pool and purchased from Twist Bioscience. The pool was amplified with KAPA polymerase (KAPA HiFi HotStart polymerase, Roche) using the following conditions: 1x KAPA HiFi 2x ReadyMix, 300 nM forward and reverse primer, 5 ng template DNA; 3 min initial denaturation at 95 °C, 20 cycles of 98 °C for 20 sec, 64 °C for 15 sec, 72 °C for 1 min and a final extension of 2 min at 72 °C. The resulting product was loaded on 2% aragose gel and gel-extracted using DNA Clean-Up & Concentration kit (Zymo Research). This pool was then cloned into a vector carrying the rest of the TadA sequence fused to dSpRYCas9 with an XTEN linker following the architecture of ABE0.1 (Supplementary sequence 1)  [27]. The library was cloned via Golden Gate assembly using BsaI restriction enzyme and T7 ligase (New England Biolabs). A total of 0.45 pmol library was mixed with 0.15 pmol vector and cycled for 5 min at 37°C and 5 min at 16°C for 30 cycles, followed by 30 min at 37°C and 5 min at 65°C. The assembled library was purified using DNA Clean-Up & Concentration kit (Zymo Research) and transformed in One Shot Top10 competent cells (Thermo Fisher Scientific) by electroporation. 950 $\mu$L SOC media were added and cells were incubated at 37°C in a shaking incubator at 200 rpm for 1 h. Samples were taken for titer plates to estimate transformation efficiency and the rest of the recovery was inoculated in 2XYT media supplemented with 36 $\mu$g/mL chloramphenicol and grown for 8h. Plasmid DNA was purified with QIAprep Spin Miniprep Kit (Qiagen) and used for selection experiments.
Base editing selection assay
To measure adenine base editing we generated a reporter carrying a carbenicillin resistance gene with a stop codon mutation (Supplementary sequence 2) and a corresponding guide RNA and kanamycin resistance gene for plasmid maintainance. The selection strain was produced by transforming Top10 E.coli cells with the reporter plasmid and preparing electrocompetent cells. Briefly, the plasmid was transformed in One Shot Top10 (Thermo Fisher Scientific) cells and a single colony was inoculated in liquid culture with 2XYT media and 50 $\mu$g/mL kanamycin. Cells were grown to mid-log phase and spun down at 3500 g, washed twice with ice-cold MiliQ water and twice with 10% glycerol (v/v). Each spin was performed for 10 min at 3500 g and after a final spin of 13 min the cell pellet was resuspended with 2 mL of 10% glycerol for 1 L of starting culture, aliquoted and flash-frozen in liquid nitrogen and stored at -80°C. For the selection experiments 100 $\mu$L of cells were transformed with 200 ng of library plasmid, electroporated and 900 $\mu$L of SOC media was added. After 1 h recovery the cells were inoculated in 4 mL 2XYT media supplemented with 50 $\mu$g/mL kanamycin (non-selective antibiotic for the reporter plasmid), 36 $\mu$g/mL chloramphenicol (for the base editor plasmid) and 0.1 % arabinose for induction of the base editor and cells were induced for 20 h (for library 1 experiments) and 16 hours (for library 2). After that cells were washed twice with 2XYT to remove inducer and split in two. Half were inoculated in 30 mL 2XYT with 50 $\mu$g/mL kanamycin and 36 $\mu$g/mL chloramphenicol (non-selective) and the other half inoculated 30 mL 2XYT with 400 $\mu$g/mL carbenicillin and 36 $\mu$g/mL chloramphenicol (selective). Cells were grown to mid to late log phase and spun down, plasmid DNA was isolated using QIAprep Spin Miniprep Kit (Qiagen). 100 ng of the plasmid DNA was used as template for a PCR1 reaction amplifying the variable base editor region and adding adapters for Illumina sequencing (Table A). The base editor portion was amplified using Q5 polymerase as follows: 100 ng plasmid DNA was used as input, 200 nM forward and reverse primer, 1x Q5 MasterMix. The reaction was carried at 98 °C for 30 sec and 20 cycles of 98 °C for 10 sec, 63 °C for 10 sec and 72 °C for 1 min, with a final extension of 5 min at 72 °C. The DNA was purified with DNA Clean-Up & Concentration kit (Zymo Research) and sent to Innovative Genomics Institute NGS core facility for PCR2 amplification and sequencing. The library was sequenced with NextSeq 1000/2000 P1 Reagents (300 Cycles) or MiSeq Reagent Kit v2 (600 cycles).
Processing sequencing data
@@
$\log e_{i}$ $=\log\left(\frac{n_{i}^{\text{s}}/{N^{\text{s}}}}{n_{i}^{\text{ns}}/N^{\text{ns}}}\right),$ (32)
@@
Pre-trained sequence generative models
@@
We trained the FMIF model using the same dataset and training protocol described in Nisonoff et al. [16] with two exceptions: (1) we used 3 encoder and decoder layers instead of 4, as both hyperparameter settings performed equally well on a held-out test set and (2) we removed all examples of engineered ABE variants from the 2021 multi-chain training dataset provided by Dauparas et al. [1]. This was done in order to ensure that FMIF does not leverage helpful information for engineering ABEs that is generally not available when engineering a new protein. Specifically, we removed PDB entry 6VPC chains E and F from the training set.
First-round library design
@@
In our first-round library, we designed around 1,000 sequences with ESM3 and FMIF, respectively, for a total library size of around 2,000 sequences. In addition, we kept three copies of the wild-type TadA sequences in the library as controls. We chose to only design the region of 86 residues on the C-terminal region of the TadA protein due to the cost constraints of gene synthesis, and we hypothesized this region might be enriched for mutations important for increasing editing activity based on prior studies [27], [81], [82]. We fixed active site residues Pro86, Cys87, and Cys90 and allowed the model to design the rest of the region. For both ESM3 and FMIF, we used 8 different values of wild-type weight $w$ linearly spaced between 0 and 3.5. We sampled from both ESM3 and FMIF using Euler integration with a timestep $\Delta t=0.002$, with a temperature of 0.5 and sampling stochasticity $\eta=20$.
Predicting ABE activity from the first-round library
The log-enrichment data collected from the first round of designs was used to design a second library enriched in highly-active base editors. Specifically, we sought to generate sequences with an average first-round log-enrichment score of at least $0.0$ (Figure 8). To do this, we trained a classifier to predict, given a sequence, whether the sequence’s log-enrichment is greater than zero or not. Guidance requires a classifier trained not only on “clean” / “un-noised” sequence-label pairs, $\{x_{1}^{i},y^{i}\}$, but also trained on “noised” sequence-label pairs $\{x_{t}^{i},y^{i}\}$, $t\in[0,1]$. This was done in two steps: first, a classifier was trained on just the “un-noised” data and then the classifier was further trained on the “noised” data.
First, the data was pre-processed by discarding sequences with highly-variable log-enrichments between the two replicates. Specifically, any sequences for which the absolute value of the difference in log-enrichments between the two replicates exceeded $0.25$ were discarded. This step filtered $114$ out of $2,000$ sequences. Next, to further mitigate the influence of experimental noise, we discarded sequences with average log-enrichment scores very close to the class boundary. Specifically, any sequences for which the average log-enrichment was between $-0.25$ and $0.25$ were discarded. This step filtered an additional $348$ sequences, resulting in a dataset of $132$ positive sequences (i.e. mean log-enrichment $\geq 0.25$) and $1,406$ negative sequences (i.e. mean log-enrichment $\leq-0.25$).

### v4

@@ 1.10 Fold class guided backbone structure generation
We first provide an overview of the experimental set-up for guiding ESM3 backbone structure generation with fold class labels from CATH [21], a hierarchical protein structure classification scheme. Concretely, we trained a fold classifier, $p_{\phi}(\mathbf{c}\,|\,\mathbf{x})$ that predict the CATH labels, $\mathbf{c}$, from the discrete structure tokens outputted by the structure encoder of ESM3, $\mathbf{x}$, on the set of CATH-labeled protein structures from the Protein Data Bank (PDB). We combined this fold class predictor with the structure generation track of ESM3, $p_{\theta}(\mathbf{x})$, to obtain a fold class conditioned backbone structure generative model, $p_{\theta,\phi}(\mathbf{x}\,|\,\mathbf{c})$ (Supplementary Figure S7a). We designed this experiment to demonstrate that ProteinGuide can be used to guide a large multimodal protein language model to generate structures.
1.10.1 Fold classifier
@@
Now we describe additional details of the fold class oracle and the fold classifier used in guidance. Since ESM3 operates on discrete structure tokens, we need a predictor which predicts the fold class label of a protein from its structure token representation. To this end, we first trained a fold class predictor that takes in discrete structure tokens as inputs. The model architecture consists of an encoder-only Transformer trunk. The final per-position embedding is mean-pooled, then fed into three parallel linear layers, each projecting the embedding to a vector with the dimension equal to the number of classes for each of the three levels of CATH hierarchies: C (“Class”), A (“Architecture”) and T (“Topology/Fold”). The C level describes the overall secondary structure-content of a structure (e.g., “mainly $\alpha$”, “mainly $\beta$”, and “mixed $\alpha/\beta$), the A level roughly clusters structures by the overall arrangement of secondary structure in space (e.g.“beta barrel”, “roll”), and the T level assigns structures by how the secondary structure elements are connected and arranged (e.g.“immunoglobulin-like”, “TIM barrel”) [21]. The model was trained on the PDB dataset following the same pre-processing procedure as in Geffner et al. [35] with a total of 214,564 structures, categorized into 5 C classes, 43 A classes, and 1,336 T classes. We fed each structure through the structure encoder of ESM3 to obtain its structure token representation. The dataset is randomly divided into training, validation, and test sets at a ratio of 8:1:1, ensuring that at least one protein from each class is included in the test set whenever possible.
@@
The model was trained with a standard cross-entropy loss applied to each of the CATH level. Some proteins have multiple CATH labels due to having multiple domains and CATH being a domain-level annotation. For such proteins, we followed Geffner et al. [35] and randomly sampled one domain label as the ground truth for each training iteration, encouraging the model to predict equal probabilities for all domain labels of a protein. During evaluation, if the model predicts any of the correct labels, the prediction is considered correct. For each protein, the loss is calculated for each level of its available CATH label. The transformer trunk of the model has a hidden dimension size of 512, 16 attention heads and 5 encoder layers. In total, the model has around 19 million parameters. The model is trained using the AdamW optimizer [63, 73] with a learning rate of 0.0001, weight decay coefficient of 0.01, a dropout rate of 0.2, a batch size of 128 and a gradient accumulation step of 4. The model is trained for around 26,000 training steps on a single NVIDIA RTX 6000 GPU. We did not systematically tune the hyperparameters of the model. On its test set, the model achieves an aggregated Micro Accuracy of 0.9678 (A level) and 0.9490 (T level) and Macro Accuracy of 0.9702 (A level) and 0.9212 (T level) [40].
@@
The fold classifier used for guidance takes in partially masked inputs. We initialized the model weights of this “noisy” classifier with those of the unnoised classifier, and trained it on the same training dataset as the unnoised classifier. We trained the noisy classifier for around 28,000 additional training steps, with the same hyperparameters as the unnoised classifier.
1.10.2 Set-up and baselines
We used ESM3 as the pre-trained model (with its structure token track) and performed fold class conditioning on the A (“Architecture”) and T (“Topology”) levels of CATH, which provide a practically informative middle ground between the coarser C level (Class), which is often too broad to be challenging or discriminative, and the finer-grained H level (Homology), which is highly specific and can be overly restrictive for controlled generation. The A level roughly clusters structures by the overall arrangement of secondary structure in space (e.g.“beta barrel”, “roll”), and the T level is a more fine-grained assignment based on how the secondary structure elements are connected and arranged (e.g.“immunoglobulin-like”, “TIM barrel”) [21]. We compared the methods on 10 most common classes and 10 randomly selected classes for both A and T level (Supplementary Table 3), evaluating on $k=100$ samples for each class. The generated structures are evaluated on two criteria: (i) whether the fold class oracle predicts the desired class as the most likely class for the generated structure, and (ii) whether the structures are considered designable, where designability is calculated following standard procedure [129]: for each generated backbone, eight sequences are sampled using ProteinMPNN [20] with a temperature of 0.1, then ESMFold [69] is used to predicted the structure of each generated sequence. The RMSD is calculated between the generated structure and the predicted structure, and a generated backbone is considered designable if the lowest RMSD—referred to as the self-consistency RMSD (scRMSD)—among the eight sampled sequences is $\leq 2$Å.
We compared ProteinGuide to unguided ESM3 sampling (with and without post hoc filtering). We additionally compared ProteinGuide to ESM3’s built-in prompt conditioning, which was trained to accept keyword-style “function tokens” as conditioning signals. Following the procedure in Hayes et al. [44] and the ESM3 GitHub repository, we converted each target CATH label to a set of ESM3-compatible keywords via an intermediate mapping through InterPro. Specifically, for Architecture (A; two-part codes) and Topology (T; three-part codes), we first mapped the CATH code to one or more associated InterPro entries, and then converted each InterPro identifier to a (possibly multi-keyword) list using ESM3’s provided InterPro-to-keyword dictionary (e.g., tokens such as beta, barrel, helix). Each keyword was provided to ESM3 as a function annotation spanning the full sequence (positions 1 through the target length); when multiple InterPro entries mapped to multiple keywords, we supplied the union of the resulting annotations.
As for the comparison to fine-tuning, we performed LoRA-based fine-tuning of ESM3 on proteins from the desired classes. Because the targets here are discrete labels, we performed fine-tuning separately for each class, substantially increasing the overall time and compute cost for fine-tuning compared to guidance. We used the same LoRA setting as previously described, with the differences that we fine-tuned the structure output head instead of the sequence output head. For each class, we performed fine-tuning on 90% of the examples, with 10% holdout for validation, training for a minimum of 500 steps and a maximum of 5000 steps, with an effective batch size of 64 and early stopping imposed if validation loss stops decreasing for more than 5 epochs. We found setting a sufficient large minimum number of training steps to be important for fine-tuning, especially for the classes with smaller number of examples. Depending on the number of samples and how quickly it converges, fine-tuning on a single class can take between several hours on the small classes to more than half a day on the larger classes on a single GPU. Due to the additional training cost of fine-tuning of this experiment, we only allowed it the same sampling budget as guidance.
@@ 1.10.2 Set-up and baselines
Following evaluation procedures in prior work [35], for each class, for every sample, we sampled its length uniformly among the lengths present for the proteins of that class in the AFDB, with a minimum length of 50 and maximum length of 275. For unguided generation, keyword prompting and fine-tuning, we used the default sampling setting of ESM3 with a sampling temperature of 0.7 and $L$ sampling steps, where $L$ is the length of the protein sequence [35]. We also experimented with using a lower sampling temperature and found that it results in ESM3 producing predominantly mainly $\alpha$ proteins. For guided generation, we used a sampling temperature of 0.3 and guidance strength $\gamma=100$. We used Euler integration with TAG with a step size $\Delta t=0.005$. We found using a sampling stochasticity $\eta=20$ to improve guided generation, but did not improve unguided generation. Unguided sampling was around 1.2 times faster than guided sampling.
1.10.3 Detailed results
Since the goal of this task is to generate more samples similar to those observed in the training distribution, and fine-tuning separately for each class allows each fine-tuned model to fully concentrate its distribution to the target class, we expect fine-tuning to be a very strong baseline. Indeed, we observed that fine-tuning to perform quite well across different levels and classes (Supplementary Figure S7b,c,d). The ProteinGuide generation also substantially enrich for the desired annotations compared to unguided generation and keyword conditioning, with success rates comparable to the strong baseline of class-specific fine-tuning (Supplementary Figure S7b,c,d). We observed keyword conditioning to be helpful for some classes but not the majority of cases. Also as expected, generation for more rare classes is more challenging (Supplementary Figure S7c,d). We also observed that across all generation methods, among the samples that were classified correctly, only a modest fraction is designable. This aligns with previous observations that ESM3 tends to produce backbones that are less designable, possibly due to the composition of its training set [35]. We expect the quality of generated structures should improve concomitantly with improved backbone generative models, independent of guidance. Examples of the generated backbone structures by guiding ESM3 with some example of CATH classes are shown (Supplementary Figure S7e).
@@ 1.10.3 Detailed results
1.11 In vivo experiment: guiding base editor for improved editing activity
@@
1.11.1 ABE Library cloning
[…]
1.11.2 Base editing selection assays
[…]
1.11.3 Processing sequencing data
@@
$\log e_{i}$ $=\log\left(\frac{n_{i}^{\text{s}}/{N^{\text{s}}}}{n_{i}^{\text{ns}}/N^{\text{ns}}}\right),$ (S35)
@@
1.11.4 Pre-trained sequence generative models
@@
We trained the FMIF model using the same dataset and training protocol described in Nisonoff et al. [82] with two exceptions: (i) we used 3 encoder and decoder layers instead of 4, as both hyperparameter settings performed equally well on a held-out test set and (ii) we removed all examples of engineered ABE variants from the 2021 multi-chain training dataset provided by Dauparas et al. [20]. This was done in order to ensure that FMIF does not leverage helpful information for engineering ABEs that is generally not available when engineering a new protein. Specifically, we removed PDB entry 6VPC chains E and F from the training set.
1.11.5 First-round library design
@@
In our first-round library, we designed around 1,000 sequences with ESM3 and FMIF, respectively, for a total library size of around 2,000 sequences. In addition, we kept three copies of the wild-type TadA sequences in the library as controls. We chose to only design the region of 86 residues on the C-terminal region of the TadA protein due to the cost constraints of gene synthesis, and we hypothesized this region might be enriched for mutations important for increasing editing activity based on prior studies [33, 91, 124]. We fixed active site residues Pro86, Cys87, and Cys90 and allowed the model to design the rest of the region. For both ESM3 and FMIF, we used 8 different values of wild-type weight $w$ linearly spaced between 0 and 3.5. We sampled from both ESM3 and FMIF using Euler integration with a timestep $\Delta t=0.002$, with a temperature of 0.5 and sampling stochasticity $\eta=20$.
1.11.6 Predicting ABE activity from the first-round library
The log-enrichment data collected from the first round of designs was used to design a second library enriched in highly-active base editors. Specifically, we sought to generate sequences with an average first-round log-enrichment score of at least $0.0$ (Supplementary Figure S8). To do this, we trained a classifier to predict, given a sequence, whether the sequence’s log-enrichment is greater than zero or not. Guidance requires a classifier trained not only on “clean” / “un-noised” sequence-label pairs, $\{x_{1}^{i},y^{i}\}$, but also trained on “noised” sequence-label pairs $\{x_{t}^{i},y^{i}\}$, $t\in[0,1]$. This was done in two steps: first, a classifier was trained on just the “un-noised” data and then the classifier was further trained on the “noised” data.
First, the data was pre-processed by discarding sequences with highly-variable log-enrichments between the two replicates. Specifically, any sequences for which the absolute value of the difference in log-enrichments between the two replicates exceeded $0.25$ were discarded. This step filtered $114$ out of $2,000$ sequences. Next, to further mitigate the influence of experimental noise, we discarded sequences with average log-enrichment scores very close to the class boundary. Specifically, any sequences for which the average log-enrichment was between 0.25-0.25 and $0.25$ were discarded. This step filtered an additional $348$ sequences, resulting in a dataset of $132$ positive sequences (i.e. mean log-enrichment $\geq 0.25$) and $1,406$ negative sequences (i.e. mean log-enrichment $\leq-0.25$).